OpenStudy (anonymous):

Write the quadratic equation given: Vertex (-2, -8) and a point: (1, 3)

OpenStudy (anonymous):

vertex is (-2,-8) you know it looks like \[y=a(x+2)^2-8\]

OpenStudy (anonymous):

you just don't know a, but you know if \[x=1\] then \[y=3\] so solve \[3=a(1+2)^2-8\] for a

OpenStudy (anonymous):

i get \[3=9a-8\] \[11=9a\] \[a=\frac{11}{9}\] hmm

OpenStudy (anonymous):

ok fine it is \[y=\frac{11}{9}(x+2)^2-8\]

OpenStudy (anonymous):

Weird... I got a=0 lol

OpenStudy (anonymous):

ohhh cuz i had it as y=a(x-x) alright thanks!